Editorial 6%, Engineering 19%. Total combined is $25,000,000. What is the annual budget.
Pleae help me. I am in the middle of a quizz and I am having trouble. Thank you

Answer by jayanthihemkos@yahoo.com(87)

Editorial 6%, Engineering 19%. Total combined is $25,000,000. What is the annual budget.
Answer:
Let the annual budget be = x
Now 6% of x +19% of x = $ 25,000,000
==> 25% 0f x = $ 25,000,000 ( add 6+19 = 25)
==> 25/100 *x = $ 25,000,000
multiply by 100/25 on both sides
==> 25/100*100/25 *x = $ 25,000,000*100/25
==> x = $25,000,000*100/25
==> x = $100,000,000
==> The annual budget = $ 100,000,000
check
6% of 100,000,000 = 6/100* 100,000,000
= 6,000,000
19% of 100,000,000 = 19/100*100,000,000
= 19,000,000
6,000,000+ 19,000,000 = 25,000,000

==> Hence the annual budget is 100,000,000